If I am going to solve a problem by completing the square, how would I know what to add to both sides?
Answer by richwmiller(17219)

After you have reduced the coefficient of x^2 to 1 then you take the coefficient of x. Now take half of it and then square it
For example
x^2+2x=8
Take the 2 of 2x and take half of it. That equals 1. Square it. That also equals 1
Add 1 to both sides.
(x+1)^2=9
Take the square root of both sides
x+1=+-3
x=-1+3=2
x=-1-3=-4
x=(2,-4)
We could have factored x^2+2x-8=0 into (x-2)*(x+4)=0